I'm having a problem with following equation using Substitution algebra. I understand substitution however this particular equation has me stumped.
4y + x = -2
x = -4y
I understand I need to plug the -4y in for the x in the first equation. What I don't understand is how to get the 4y on the other side of = sign in order to isolate the y.
Thanks.

Answer by Alan3354(69443)

4y + x = -2
x = -4y
---------------
If you sub for x:
4y + (-4y) = -2
0 = -2
There's no solution.
-----------------
4y + x = -2 --> x = -4y - 2
x = -4y
If you graph these, you get 2 parallel lines. No intersection, no solution.